# ICE: Intrinsic Concept Extraction from a Single Image via Diffusion Models

### v1

@@ Abstract
The inherent ambiguity in defining visual concepts poses significant challenges for modern generative models, such as the diffusion-based Text-to-Image (T2I) models, in accurately learning concepts from a single image. Existing methods lack a systematic way to reliably extract the interpretable underlying intrinsic concepts. To address this challenge, we present ICE, short for Intrinsic Concept Extraction, a novel framework that exclusively utilizes a T2I model to automatically and systematically extract intrinsic concepts from a single image. ICE consists of two pivotal stages. In the first stage, ICE devises an automatic concept localization module to pinpoint relevant text-based concepts and their corresponding masks within the image. This critical stage streamlines concept initialization and provides precise guidance for subsequent analysis. The second stage delves deeper into each identified mask, decomposing the object-level concepts into intrinsic concepts and general concepts. This decomposition allows for a more granular and interpretable breakdown of visual elements. Our framework demonstrates superior performance on intrinsic concept extraction from a single image in an unsupervised manner. Project page: https://visual-ai.github.io/ice
Figure 1: We showcase a structured approach for defining visual concepts within an image, where object-level concepts are identified and analyzed to reveal their underlying intrinsic attributes, such as color, material, and shape. We present the ICE (Intrinsic Concept Extraction) framework, which leverages Text-to-Image (T2I) models to systematically discover these concepts, providing a more effective method for learning visual concepts.
@@ 1 Introduction
Addressing these challenges requires a fundamental re-evaluation of how we define and learn visual concepts within an image. Previous works, such as Inspiration Tree [34], have attempted to construct hierarchical representations by organizing tokens into tree structures to capture different concepts within an image. Similarly, LangInt [17] uses Visual Question Answering (VQA) techniques to guide concept learning, but it requires training separate encoders for each concept axis, thereby failing to fully utilize their inherent hierarchical structures. Consequently, there is a need for more systematic and structured methods to define and learn concepts from images, enabling a more granular understanding of the fundamental building blocks of visual representation.
@@ 2 Related Work
Figure 2: Concept definition hierarchy illustrating how object-level concepts are decomposed into intrinsic attributes, including object type, colour, material and other intrinsics.
@@ 4.1 Stage One: Automatic Concept Localization
where $\mathbf{m}_{i}$ delineates the region of $\mathbf{x}$ associated with the concept $c_{i}$. In our implementation, without introducing any external pretrained model, we utilize a training-free image segmentor that leverages the self-attention layers within the pre-trained Stable Diffusion model to produce high-quality segmentation masks without any additional training.
@@ 4.2 Stage Two: Structured Concept Learning
Phase two: learning intrinsic concepts. The second phase delves into decomposing the learned object-level concepts into intrinsic attributes. For each intrinsic concept (e.g., colour, material), we introduce corresponding intrinsic tokens $c_{j}^{\text{intrinsic}}$ associated with object-level concepts $c_{i}$. The intrinsic concept tokens are anchored using combined representations of $c_{i}^{\text{conspec}}$ and $c_{i}^{\text{inspec}}$. Specifically, for an intrinsic concept $j$, the anchor text $\texttt{anchor}_{j}$ is formulated as:
@@
where $\texttt{intrinsic}_{j}$ denotes the specific attribute (e.g., color). The triplet loss $\mathcal{L}_{\text{triplet}}^{\text{intrinsic}}$ for intrinsic concept $j$ is then defined as:
@@ 4.3 Concept refinement
Optimizing intrinsic concepts solely through token embeddings is challenging, often leading to misaligned and inconsistent representations of specific visual attributes such as color and texture. To enhance the accuracy and consistency of these intrinsic concepts, after completing Phase two, we perform concept refinement by fine-tuning both the U-Net denoising network and the text encoder $\mathcal{E}$ for a limited number of training steps. This fine-tuning ensures that the intrinsic concepts are precisely aligned with their corresponding visual attributes.
@@ 5.2 Qualitative results
First, given an input image $x$, ICE extracts corresponding concept masks and retrieves relevant text-based descriptors, as shown in the second column of the figure. These masks precisely delineate regions within the image, ensuring accurate localization of each concept. Next, leveraging the extracted masks and retrieved texts, ICE learns distinct object-level concepts, displayed in the third column. For example, the framework identifies objects such as “Bust” and “Cone” and associates them with their respective semantic descriptors. Finally, utilizing the information from the object-level concepts, ICE systematically discovers intrinsic concepts, including attributes like material and colour, through its second-stage processing. This decomposition enables the framework to uncover compositional intrinsic concepts such as the object type, material, and colour, as demonstrated in the fourth column.
@@ 6 Conclusion
Since our proposed ICE framework introduces a novel approach to extracting intrinsic concepts, we design a comprehensive evaluation method to assess the quality of these intrinsic concepts. We call this Intrinsic Concept benchmark (ICBench). To achieve this, we expand upon the dataset utilised by the Unsupervised Concept Extraction (UCE) benchmarks [11], referred to as the D1 dataset (96 images), which is based on Unsplash  [^2] https://unsplash.com/ images. For each extracted concept mask in every image, we provide detailed concept descriptions, denoted as $description_{j}$, encompassing $intrinsic_{j}$ attributes such as object, material, and colour. These descriptions, as illustrated in Figure A, are generated using GPT-4o model [14] to ensure accurate and reliable descriptions.
Figure A: Representative intrinsic concept evaluation descriptions generated by GPT-4o model for each segmentation mask produced by the ICE framework. Each description includes detailed attributes such as object (Obj), material (Mat), and colour (Col).
@@
In this section, we present additional qualitative results generated by our ICE framework to demonstrate its effectiveness in accurately localising and decomposing visual concepts. Figure B showcases a variety of unlabelled images along with the corresponding extracted concepts and their segmentation masks. These examples illustrate the framework’s capability to identify and segment distinct objects within an image, as well as decompose them into their intrinsic attributes such as object, colour, and material. The results highlight the precision and interpretability of the concept representations produced by ICE, reinforcing the advantages of our structured approach to learning concepts.
@@
Our ICE framework addresses this limitation by decomposing object-level concepts into their intrinsic attributes through Stage Two: Structured Concept Learning. This decomposition facilitates the disentanglement of attributes like object, material, and colour, allowing for flexible manipulation of each attribute without altering the overall object identity. Figure C illustrates an example where ICE enables compositional generation of an object’s colour and material independently. By using ICE framework, we can specifically edit or alter the material while keeping the colour unchanged. This granular control over intrinsic attributes empowers users to perform precise and targeted compositional concept edits, enhancing the flexibility and utility of image generation tasks. By enabling the independent manipulation of intrinsic attributes, ICE not only enhances the customization capabilities of image generation but also paves the way for more sophisticated and user-driven image generation applications. This advancement overcomes the limitations of current object-level concept learning methods, providing a more nuanced and flexible approach to text-to-image generation.
@@
Our ICE framework is divided into two stages: automatic concept localisation (Stage One) and structured concept learning (Stage Two). In this section, we showcase the potential of our automatic concept localisation module for zero-shot unsupervised segmentation. Specifically, our Stage One module utilises DiffSeg [32] as the zero-shot segmentor $\mathcal{S}$, enabling segmentation without any additional training or reliance on language dependencies. DiffSeg leverages a pre-trained Text-to-Image (T2I) Stable Diffusion model to perform zero-shot segmentation tasks. Specifically, it employs an iterative merging process based on measuring the distribution among self-attention maps extracted from the T2I model. We compare Stage One of our method with DiffSeg and some traditional clustering techniques such as $k$-means-S [32], which utilises $k$-means clustering with a predefined number of clusters based on the ground truth classes, and the DBSCAN [8] algorithm, known for its density-based clustering approach. These comparisons provide a comprehensive evaluation of our ICE framework’s performance in unsupervised segmentation tasks.
@@
The development of unsupervised compositional concept discovery frameworks like ICE has significant implications for the field of computer vision and beyond. By enabling more granular and interpretable concept learning, ICE enhances the capabilities of generative AI applications in tasks such as image content personalization. These advancements can lead to more intuitive and user-friendly applications, such as advanced image generation tools and more accurate compositional concept generation. However, the use of pre-trained T2I diffusion models in ICE implies reliance on datasets that may contain inherent biases. These biases can be inadvertently perpetuated or even amplified in the generated concepts, leading to unfair or discriminatory outcomes. It is imperative to address these biases through careful dataset curation and model auditing to ensure the equitable and responsible deployment of ICE in real-world applications.

### v2

@@ Abstract
The inherent ambiguity in defining visual concepts poses significant challenges for modern generative models, such as the diffusion-based Text-to-Image (T2I) models, in accurately learning concepts from a single image. Existing methods lack a systematic way to reliably extract the interpretable underlying intrinsic concepts. To address this challenge, we present ICE, short for Intrinsic Concept Extraction, a novel framework that exclusively utilises a T2I model to automatically and systematically extract intrinsic concepts from a single image. ICE consists of two pivotal stages. In the first stage, ICE devises an automatic concept localization module to pinpoint relevant text-based concepts and their corresponding masks within the image. This critical stage streamlines concept initialization and provides precise guidance for subsequent analysis. The second stage delves deeper into each identified mask, decomposing the object-level concepts into intrinsic concepts and general concepts. This decomposition allows for a more granular and interpretable breakdown of visual elements. Our framework demonstrates superior performance on intrinsic concept extraction from a single image in an unsupervised manner. Project page: https://visual-ai.github.io/ice
Figure 1: We showcase a structured approach for defining visual concepts within an image, where object-level concepts are identified and analyzed to reveal their underlying intrinsic attributes, such as object category, colour, and material. We present the ICE (Intrinsic Concept Extraction) framework, which leverages Text-to-Image (T2I) models to systematically discover these concepts, providing a more effective method for learning visual concepts.
@@ 1 Introduction
Addressing these challenges requires a fundamental re-evaluation of how we define and learn visual concepts within an image. Previous works, such as Inspiration Tree [34], have attempted to construct hierarchical representations by organizing tokens into tree structures to capture different concepts within an image. Similarly, LangInt [17] utilises Visual Question Answering (VQA) techniques to guide concept learning, but it requires training separate encoders for each concept axis, thereby failing to fully utilize their inherent hierarchical structures. Consequently, there is a need for more systematic and structured methods to define and learn concepts from images, enabling a more granular understanding of the fundamental building blocks of visual representation.
@@ 2 Related Work
Figure 2: Concept definition hierarchy illustrating how object-level concepts are decomposed into intrinsic attributes, including object category type, colour, material and other intrinsics.
@@ 4.1 Stage One: Automatic Concept Localization
where $\mathbf{m}_{i}$ delineates the region of $\mathbf{x}$ associated with the concept $c_{i}$. In our implementation, without introducing any external pretrained model, we utilize a training-free image segmentor that leverages the self-attention layers within the pretrained Stable Diffusion model to produce high-quality segmentation masks without any additional training.
@@ 4.2 Stage Two: Structured Concept Learning
Phase two: learning intrinsic concepts. From the previous phase, we obtained the “object category" intrinsic concept by learning $c_{i}^{\text{conspec}}$. The second phase delves into decomposing the learned object-level concepts into other intrinsic attributes (e.g., colour, material). For each intrinsic concept, we introduce corresponding intrinsic tokens $c_{j}^{\text{intrinsic}}$ associated with object-level concepts $c_{i}$. The intrinsic concept tokens are anchored using a text prompt, tailored to each intrinsic concept. Specifically, for an intrinsic concept $j$, the anchor text $\texttt{anchor}_{j}$ is formulated as:
@@
where $\texttt{intrinsic}_{j}$ denotes the specific attribute (e.g., colour). The triplet loss $\mathcal{L}_{\text{triplet}}^{\text{intrinsic}}$ for intrinsic concept $j$ is then defined as:
@@ 4.3 Concept refinement
Optimizing intrinsic concepts solely through token embeddings is challenging, often leading to misaligned and inconsistent representations of specific visual attributes such as colour and texture. To enhance the accuracy and consistency of these intrinsic concepts, after completing Phase two, we perform concept refinement by fine-tuning both the U-Net denoising network and the text encoder $\mathcal{E}$ for a limited number of training steps. This fine-tuning ensures that the intrinsic concepts are precisely aligned with their corresponding visual attributes.
@@ 5.2 Qualitative results
First, given an input image $x$, ICE extracts corresponding concept masks and retrieves relevant text-based descriptors, as shown in the second column of the figure. These masks precisely delineate regions within the image, ensuring accurate localization of each concept. Next, leveraging the extracted masks and retrieved texts, ICE learns distinct object-level concepts, displayed in the third column. For example, the framework identifies objects such as “Bust” and “Cone” and associates them with their respective semantic descriptors. Finally, utilizing the information from the object-level concepts, ICE systematically discovers intrinsic concepts, including attributes like material and colour, through its second-stage processing. This decomposition enables the framework to uncover compositional intrinsic concepts such as the object category type, material, and colour, as demonstrated in the fourth column.
@@ 6 Conclusion
Since our proposed ICE framework introduces a novel approach to extracting intrinsic concepts, we design a comprehensive evaluation method to assess the quality of these intrinsic concepts. We call this Intrinsic Concept benchmark (ICBench). To achieve this, we expand upon the dataset utilised by the Unsupervised Concept Extraction (UCE) benchmarks [11], referred to as the D1 dataset (96 images), which is based on Unsplash  [^2] https://unsplash.com/ images. For each extracted concept mask in every image, we provide detailed concept descriptions, denoted as $description_{j}$, encompassing $intrinsic_{j}$ attributes such as object category, material, and colour. These descriptions, as illustrated in Figure A, are generated using GPT-4o model [14] to ensure accurate and reliable descriptions.
Figure A: Representative intrinsic concept evaluation descriptions generated by GPT-4o model for each segmentation mask produced by the ICE framework. Each description includes detailed attributes such as object category (Obj), material (Mat), and colour (Col).
@@
In this section, we present additional qualitative results generated by our ICE framework to demonstrate its effectiveness in accurately localising and decomposing visual concepts. Figure B showcases a variety of unlabelled images along with the corresponding extracted concepts and their segmentation masks. These examples illustrate the framework’s capability to identify and segment distinct objects within an image, as well as decompose them into their intrinsic attributes such as object category, colour, and material. The results highlight the precision and interpretability of the concept representations produced by ICE, reinforcing the advantages of our structured approach to learning concepts.
@@
Our ICE framework addresses this limitation by decomposing object-level concepts into their intrinsic attributes through Stage Two: Structured Concept Learning. This decomposition facilitates the disentanglement of attributes like object category, material, and colour, allowing for flexible manipulation of each attribute without altering the overall object identity. Figure C illustrates an example where ICE enables compositional generation of an object’s colour and material independently. By using ICE framework, we can specifically edit or alter the material while keeping the colour unchanged. This granular control over intrinsic attributes empowers users to perform precise and targeted compositional concept edits, enhancing the flexibility and utility of image generation tasks. By enabling the independent manipulation of intrinsic attributes, ICE not only enhances the customization capabilities of image generation but also paves the way for more sophisticated and user-driven image generation applications. This advancement overcomes the limitations of current object-level concept learning methods, providing a more nuanced and flexible approach to text-to-image generation.
@@
Our ICE framework is divided into two stages: automatic concept localisation (Stage One) and structured concept learning (Stage Two). In this section, we showcase the potential of our automatic concept localisation module for zero-shot unsupervised segmentation. Specifically, our Stage One module utilises DiffSeg [32] as the zero-shot segmentor $\mathcal{S}$, enabling segmentation without any additional training or reliance on language dependencies. DiffSeg leverages a pretrained Text-to-Image (T2I) Stable Diffusion model to perform zero-shot segmentation tasks. Specifically, it employs an iterative merging process based on measuring the distribution among self-attention maps extracted from the T2I model. We compare Stage One of our method with DiffSeg and some traditional clustering techniques such as $k$-means-S [32], which utilises $k$-means clustering with a predefined number of clusters based on the ground truth classes, and the DBSCAN [8] algorithm, known for its density-based clustering approach. These comparisons provide a comprehensive evaluation of our ICE framework’s performance in unsupervised segmentation tasks.
@@
The development of unsupervised compositional concept discovery frameworks like ICE has significant implications for the field of computer vision and beyond. By enabling more granular and interpretable concept learning, ICE enhances the capabilities of generative AI applications in tasks such as image content personalization. These advancements can lead to more intuitive and user-friendly applications, such as advanced image generation tools and more accurate compositional concept generation. However, the use of pretrained T2I diffusion models in ICE implies reliance on datasets that may contain inherent biases. These biases can be inadvertently perpetuated or even amplified in the generated concepts, leading to unfair or discriminatory outcomes. It is imperative to address these biases through careful dataset curation and model auditing to ensure the equitable and responsible deployment of ICE in real-world applications.